The owner of ABC Co. suspects abnormalities in the processing of accounts receivable. This is ABC Co.'s first year of business, and the following information is available.

Collections from customers, $203,000

Merchandise purchased, $325,000

Ending merchandise inventory, $95,000

Goods are marked to sell at 50% above cost.

Accounts receivable balance per the ledger, $87,000

Which of the following statements is correct?

Select one:

a. The owner is correct; accounts receivable is understated in the ledger by $27,000.

b. $87,000 is the appropriate balance; the ledger is correct.

c. The owner is correct; accounts receivable is understated in the ledger by $55,000.

d. The owner is correct; accounts receivable is understated in the ledger by $142,000.
Answer #1

Cost of goods sold = (325000 - 95000) = $230,000

Sale 230000 *1.50= $345,000

Actual accounts receivable balance = Sale - Collections from customers

= 345000 - 203000 = $142000

accounts receivable is understated in the ledger by = 142000 - 87000 = $55000

option c. is correct answer.